# SoK: Membership Inference Attacks on LLMs are Rushing Nowhere (and How to Fix It)

### v2

@@ 5.3 Overlap between members and non-members
This inherent overlap of text is a well-known issue which has been heavily studied in the context of properly evaluating newly released LLMs. For instance, in GPT-3 [10], authors found a significant overlap between the training data and established evaluation benchmarks. To evaluate how the model generalizes to unseen data, they aggressively deduplicate ($13$-gram overlaps) between the training data and benchmarks. This methodology (decontamination) has since been adopted and further refined by subsequent model developers [99, 26, 95]. Even these ‘string matching’ methods however remain insufficient to address near-duplicates, semantic duplicates or translations [102].
@@
The overlap also poses challenges for MIAs. Indeed, non-member data that is partially seen during training may act as quasi-members, complicating the inference process. Moreover, recent research suggests that near-duplicates in training fuels memorization. Shilov et al. [87] demonstrate that the inclusion of near-duplicates in the training dataset of an LLM enhances the distinguishability of a sequence in MIAs. Further, it is well established that deduplication, either literal [52] or semantic [1], improves LLM performance, suggesting that near-duplicates contribute to redundant memorization.
@@
Recently, Duan et al. [27] applied common techniques used for decontamination on a randomized split of train and test data from the Pile [31]. These splits have been used to train and validate some open-source models [7, 8], which recent research [55, 67, 27] suggests to offer a randomized privacy game to evaluate MIAs against LLMs. Duan et al [27] finds that these splits, just as encountered for evaluation benchmarks, can indeed have substantial overlap, making it harder for MIAs to succeed. They report an average of 32.5% $7$-gram overlap between members and non-members for Wikipedia articles from the Pile [31], even reaching up to 76.9% overlap for data collected from Github.
6 Towards proper MIA evaluation
@@
Aware of the importance of, and challenges associated with, creating a randomized setup to evaluate MIAs against LLMs, we here review and discuss potential solutions proposed in the literature.
6.1 Randomized test split
@@
Table 3: AUC (mean±$\pm$±standard deviation) for the bag of words classifier applied to subsets of The Pile, either directly sampled from its uncopryighted version [2] or from MIMIR [27] (which applied further deduplication).
The privacy game for a fixed model [104] only requires one training run, and could be naturally instantiated alongside regular training, making it feasible even for expensive LLMs.
Some pretrained LLMs make their entire training dataset open-source, including a randomized and controlled split of ’test’ data. While this set of test documents is primarily intended for a fair evaluation of the model’s ability to generalize its language modeling capabilities to unseen language, such test sets might prove very valuable to study and evaluate MIAs [27, 55, 67]. The training and validation splits are for instance made available for the suite of Pythia models [7], GPT-NEO [8], the OLMo model family [34] and the CroissantLLM project [30]. Finally, some initiatives such as OpenLLaMA [32] and LLM360 [61] release the exact training data sequences used, making it possible to know whether particular samples have been seen by intermediate model checkpoints or not.
First, we confirm whether the distribution shift identified earlier for members and non-members collected post-hoc, is now eliminated in a real-world, randomized setup. We consider three datasets. First, we consider $1,000$ members and non-members directly sampled from the train and test splits of (uncopyrighted version of) The Pile [2] (used to train and validate Pythia [7] and GPT-NEO [8]). Next, we consider both versions of the MIMIR dataset (13_0.8 and 7_0.2) [27]. As discussed in Sec. 5.3, Duan et al. [27] identified a substantial overlap between members and non-members in the randomized split from The Pile, making the notion of membership poorly defined and making it harder for MIAs to succeed. To mitigate this issue, they apply a deduplication strategy on the non-member data, removing non-members with more than 80% overlap in $13$-grams (13_0.8) and more aggressively, samples with more than 20% overlap in $7$-grams (7_0.2). Authors release the result as the MIMIR dataset. We train and evaluate a bag of words classifier on all three datasets and report the results in Table 3.
@@
First, we confirm that, for the data directly sampled from the train and test split across all $7$ subsets from the Pile (left column), the bag of words classifier barely performs better than a random guess baseline; consistent with what we expect from a randomized split.
@@
For the MIMIR dataset [27], however, the results (right columns) are more nuanced. For the moderate level of deduplication (13_0.8) for almost all data subsets, we find the AUC to remain close to a random guess baseline. This means that the distribution shift remains mitigated, while removing non-members with substantial overlap in the training dataset. For data collected from Github, however, the AUC of our bag of words classifier now corresponds to $0.68$, suggesting that there is a meaningful difference between members and non-members. From the analysis in Duan et al. [27], this suggests that the overlap between non-members and members in Github is so substantial that the non-members remaining after deduplication are mostly outliers, which initiates a new distribution shift. For the more aggressive deduplication (7_0.2), we find similarly large AUC for almost all subsets. This suggests that deduplicating too aggressively post-hoc might distort the exact properties of a randomized setup that are desired to evaluate MIAs. Future research is needed to understand the level of deduplication which can be applied post-hoc to mitigate the effect of the overlap issue, while not introducing any distribution shift.
Apart from the Github dataset split and the aggressive deduplication (7_0.2), the bag of words AUC suggests that the MIMIR dataset allows for a randomized setup to evaluate MIAs against LLMs. We instantiate a range of MIAs from the literature on the MIMIR dataset [27] (13_0.8, and not considering the Github subset) using Pythia-dedup-6.9B [7]. Table 4 summarizes the MIA AUC for this setup. We confirm, across a range of novel MIA methods proposed, the attack performance remains only barely better than random. Importantly, also the attacks proposed most recently [109, 101, 97, 105, 108, 45], which claim to out-perform prior methods in flawed evaluation setups, do not perform any better than a random guess in this clean setup.
@@
We also instantiate document-level MIAs on a randomized test split of entire ArXiv documents from the Pile [2]. We consider both sequence-level MIAs from the literature, scaled to the document-level as proposed by Shi et al. [86] and formalized in Sec. 5.2 as well as the document-level MIA proposed by Meeus et al. [69]. Also for document-level MIAs, we find the MIA AUC to barely perform any better than random. We provide the results in Table 8 with the detailed experimental setup in (Appendix B).
@@
Table 4: Benchmarking MIAs on the MIMIR dataset [27] and Pythia-dedup-6.9B [7]. Sequences are sampled from the train (members) and test (non-members) split of The Pile [31]. MIA AUC computed using 500 members and 500 non-members.
@@
Limitations. While held-out test splits generally provide a randomized setup for privacy games, the substantial overlap between members and non-members is a core limitation. Post-hoc deduplication could potentially mitigate the overlap, yet determining the appropriate level of deduplication is difficult to evaluate and might distort the randomization. Its effectiveness likely varies depending on the dataset type, as demonstrated in the case of the GitHub subset.
@@
The practical significance of the overlap issue is a nuanced, perhaps philosophical, question. To rigorously examine what an LLM memorizes and how effectively MIAs can distinguish between members and non-members, one might prefer a setup with minimal overlap. Yet, in real-world applications, overlap is an inherent feature of text. To accurately infer whether a specific piece of text, such as a copyright-protected book or a medical record, was used to train an LLM, it is likely crucial to consider any overlap with other member data. For example, well-known quotes from a book might be widespread online (for which copyright might not apply), while generic elements of a medical record, even if present in the training data, may not pose a privacy risk. Whether MIAs -to e.g. study privacy risk in practice- should be evaluated on datasets with an overlap thus remains an open question.
6.2 Injection of randomized unique sequences
@@
Table 5: Benchmarking MIAs across copyright traps for CroissantLLM [30]: MIA AUC for $500$ members and non-members.
Instead of relying on a randomized test split of documents, prior work has proposed the injection of highly unique sequences, mitigating the overlap, to evaluate sequence-level MIAs.
For smaller language models, the controlled injection of specific sequences in the training data has been studied extensively. Henderson et al. [36] injects a variety of key-value pairs in a sequence-to-sequence dialogue model, while Carlini et al. [15] proposes the injection of canaries, unique sequences of random numbers, in the training set of LSTM models; both to measure how these sequences are memorized. Further work has, for instance, injected unique sequences to study memorization of federated learning [92] or of word embeddings [93].
More recently, Wei et al. [100] investigates how data watermarks, i.e. random sequences of tokens injected into the training dataset, are memorized by medium scale language models (70M parameters trained on 100M tokens). Further, Meeus et al. [70] injected canaries (which they call traps sequences) into the pretraining dataset of CroissantLLM [30], a 1.3B parameter model trained on 3T tokens. The canaries are synthetically generated, unique sequences, injected with varying sequence length, number of repetitions, and perplexity. As the canaries are synthetically generated (by sampling from a pretrained LLM), the same process allows generating new, unique canaries coming from the exact same distribution, yet not injected into the training dataset of CroissantLLM [30] (non-members).
While Meeus et al. [70] use this setup to study how the inclusion of canaries enables content owners to track whether their content has been used for LLM training, the setup also provides a randomized and controlled setup to evaluate MIAs against LLMs. We confirm this to be the case by fitting the bag-of-words classifier on copyright traps (using $500$ members and non-members), recovering an AUC of $0.543\pm 0.037$ for sequences of $100$ tokens. This confirms it to be a randomized setup for evaluating sequence-level MIAs.
Compared to the randomized setup available with a randomized test split (Sec. 6.1), the use of randomized unique sequences to evaluate MIAs might have two distinct benefits.
First, when the randomized sequences are highly unique, they could effectively eliminate the challenge of overlap between members and non-members. Meeus et al. [70] generate synthetic sequences of high perplexity, suggesting that these sequences would likely be more unique compared to the rest of the target model training data than a randomized test split.
Further, injected sequences can be repeated. As it is well established that repeated sequences are memorized more [13, 44], injecting with a varying number of repetitions would provide an MIA evaluation setup across multiple regimes of LLM memorization. This could also enable the comparison of different MIAs with (more) statistical significance.
We instantiate MIAs from the literature on the dataset of member and non-member canaries as made available by Meeus et al. [70] on the open-source CroissantLLM [30]. Table 5 summarizes the MIA AUC across a number of repetitions. For $1,000$ repetitions, we recover MIAs reaching a performance significantly better than a random guess, with Ratio-LLaMA-2 [16] (AUC of $0.70$) and Min-KProb [86] (AUC of $0.63$) the best performing methods.
Limitations. While generally a great randomized setup to develop and evaluate MIAs against LLMs, using the randomized unique sequences to evaluate MIAs also comes with its limitations.
@@
First, the MIA performance might only represent a worst-case risk and be not representative of real-world data. Indeed, highly unique sequences mitigate any overlap between members and non-members, making it easier for MIAs to succeed, especially when repeated during training. Moreover, Meeus et al. [70] shows that sequences with larger perplexity are more vulnerable, and considers sequences with a perplexity higher than naturally occurring ones. Hence, evaluating MIAs on unique sequences might give a worst-case MIA risk compared to real-world data. While this might be beneficial for some use-cases (e.g. privacy auditing of methods with formal guarantees [63, 64, 91], detecting copyright-protected content [70, 100]), evaluating MIAs on (artificially) unique sequences might not be representative of risks in practice.
@@
Second, the success of MIAs might depend on certain properties of the sequence. For instance, the Ratio MIA [16] relies on the use of a reference model, which should normalize the membership score by how ’complex’ a certain sequence is compared to other natural language. Surely, the extent to which this normalization is effective for an MIA against a target LLM depends on the choice of reference model, its training data and how well it generalizes to unseen data. Meeus et al. [70] use LLaMA-2 7B [95] to synthetically generate the trap sequences, which likely affects the success of certain reference models in detecting such traps. Also Duan et al. [16] finds "choosing the right reference model for a target LLM challenging and largely empirical". Moreover, we hypothesize that certain MIAs may vary in effectiveness based on sequence length (e.g. Min-KProb [86], which relies on a subset of token-level probabilities). Similarly, some MIAs (e.g. SMIA [73]) depend on the semantic meaning of a target sequence, which might be less effective for sequences with very high perplexity due to the lack of semantically meaningful embeddings in such cases. This suggests that using randomized unique sequences as an MIA evaluation method could influence the relative performance of MIA methodologies.
6.3 Finetuning
@@
Leveraging finetuning for MIA evaluation has distinct benefits compared to previous approaches. First, similar to the injection of unique sequences with varying repetitions, finetuning could be done for varying memorization regimes (adjusting e.g. number of parameters updated during finetuning, dataset size, learning rate [71]). As such, this setup would be well suited to compare MIA performance across methods.
Further, from a practical perspective, it is very sensible to study MIAs in post-pretraining attack models. Finetuning (or continued pretraining) is often done on smaller, domain-specific, high quality datasets which are often more susceptible to contain sensitive information. It is cost effective to start off from generalist model checkpoints (e.g. LLaMA-3 [3], Mistral-7B [43]) and train on a domain-specific, causal language modeling task for a few billions of tokens to specialize models. Examples include finetuning for legal domain adaptation [19], for multilingual abilities [4], and for medical domain adapation [49]. Moreover, annealing, or post-training techniques in which LLMs are trained on high-quality data splits after their initial pretraining phase, is now understood to be essential to boost model performance [38, 3]. Such high-quality data (news articles, textbooks, synthetic data) is often the category most likely to be copyright-protected.
Limitations. For some MIA use-cases, including detecting the use of copyright-protected content or evaluation benchmarks during pretraining, the evaluation of MIAs remains generally desired for pretrained models rather than for finetuned models. How MIA evaluation on finetuned models translates to larger models, datasets or training hyperparameters remains an open question [6].
6.4 Post-hoc control methods

### v3

@@ V-C Overlap between members and non-members
This inherent overlap of text is a well-known issue which has been heavily studied in the context of properly evaluating newly released LLMs. For instance, in GPT-3 [9], authors found a significant overlap between the training data and established evaluation benchmarks. To evaluate how the model generalizes to unseen data, they aggressively deduplicate ($13$-gram overlaps) between the training data and benchmarks. This methodology (decontamination) has since been adopted and further refined by subsequent model developers [98, 26, 94]. However, even these exact deduplication methods remain insufficient to address near-duplicates, semantic duplicates or translations [101], highlighting the persistent challenge of effective decontamination.
@@
The overlap also poses similar challenges for MIAs. Indeed, non-member data that is partially seen during training may act as quasi-members, potentially making the membership inference task poorly defined. Moreover, recent research suggests that the inclusion of near-duplicates of members in the training data enhances the members’ distinguishability in MIAs [86], further complicating the inference process.
@@
Recent works [54, 27, 65] have proposed to use the randomized split of train and test data from the Pile [31] to evaluate MIAs against LLMs. Indeed, these splits have been used to train and validate some open-source models [6, 7] and thus enable to instantiate a proper, fixed-model privacy game. However, Duan et al. [27] finds that these splits, just as encountered for evaluation benchmarks, can indeed have substantial overlap, making it harder for MIAs to succeed. They report an average of 32.5% $7$-gram overlap between members and non-members for Wikipedia articles from the Pile [31], even reaching up to 76.9% overlap for data collected from Github.
VI Towards proper MIA evaluation
@@
Recognizing the importance and challenges of creating a randomized setup for evaluating MIAs against LLMs, we review and discuss potential solutions from the literature.
VI-A Randomized test split
The privacy game for a fixed model [103] only requires one training run, and could be naturally instantiated alongside regular training. When planned and established before training begins, this remains feasible even for expensive LLMs.
Some pretrained LLMs make their entire training dataset open-source, including a randomized and controlled split of ‘test’ data. While this set of test documents is primarily intended for a fair evaluation of the model’s ability to generalize its language modeling capabilities to unseen language, such test sets might prove very valuable to study and evaluate MIAs [27, 54, 65]. The training and validation splits have been, for instance, made available for models such as Pythia [6], GPT-NEO [7], OLMo [34] and CroissantLLM [30].
First, we use our bag of words classifier to analyze whether the distribution shift identified earlier for members and non-members collected post-hoc, is now eliminated in a real-world, randomized set of members and non-members. We consider three such datasets. First, we consider $1,000$ members and non-members directly sampled from the train and test splits of (the uncopyrighted version of) The Pile [1] (used to train and validate Pythia [6] and GPT-NEO [7]). Next, we consider both versions of the MIMIR dataset (13_0.8 and 7_0.2) [27]. As discussed in Sec. V-C, Duan et al. [27] identified a substantial overlap between members and non-members in the randomized split from The Pile, making the notion of membership poorly defined and making it harder for MIAs to succeed. To mitigate this issue, they apply a deduplication strategy on the non-member data, removing non-members with more than 80% overlap in $13$-grams (13_0.8) and more aggressively, samples with more than 20% overlap in $7$-grams (7_0.2). Authors release the result as the MIMIR dataset.
TABLE III: AUC (mean±$\pm$±standard deviation) for the bag of words classifier applied to subsets of The Pile, either directly sampled from its uncopryighted version [1] or from MIMIR [27] (which applied further deduplication).
@@ VI-A Randomized test split
The bag of words performance in Table III (first column) first shows that, for the data directly sampled from the train and test split across all $7$ subsets from the Pile, the bag of words classifier barely performs better than a random guess; consistent with what we expect from a randomized split.
@@
For the MIMIR dataset [27], however, the results (second and third columns) are more nuanced. For the moderate level of deduplication (13_0.8), we find the AUC to remain close to a random guess baseline for almost all data subsets. This means that the deduplication did not introduce a distribution shift detected by the bag of words classifier, while removing non-members with substantial overlap in the training dataset. For data collected from Github, however, our bag of words classifier obtains an AUC of $0.68$, suggesting that there is a meaningful distribution shift between members and non-members. This implies that the overlap between members and non-members for data collected from Github is so substantial that the non-members remaining after deduplication are mostly outliers, which initiates the distribution shift. Duan et al. [27] indeed confirms that the overlap is significantly more substantial for Github than for other data subsets and illustrates specific outliers that remain after deduplication. For the more aggressive deduplication (7_0.2), we find similarly large AUC for almost all subsets. This suggests that deduplicating too aggressively post-hoc might distort the exact properties of a randomized setup that are desired to evaluate MIAs. Future research is needed to understand the level of deduplication which can be applied post-hoc to mitigate the effect of the overlap issue, while not introducing any distribution shift. Our results suggest that MIMIR (13_0.8), apart from the data from Github, would serve as a randomized setup to evaluate MIAs against LLMs. However, we note again that the AUC achieved by the bag of words classifier only represents a lower bound; if its performance is close to a random guess, this does not necessarily imply the absence of any distribution shift.
We instantiate a range of MIAs from the literature on the MIMIR dataset [27] (13_0.8, and not considering the Github subset) using Pythia-dedup-6.9B [6]. Table IV shows that the original attacks [105, 15, 66, 85] do not perform significantly better than random, and neither do any of the $5$ recently proposed attacks [109, 96, 104, 107, 45] which allegedly outperformed the original ones.
@@
We also instantiate document-level MIAs on a randomized split of entire ArXiv documents from the Pile [1]. We consider both MIAs applied to all sequences in the document, scaled to the document-level as proposed by Shi et al. [85] (formalized in Sec. V-B) as well as the document-level MIA (using a meta-classifier) proposed by Meeus et al. [68]. Also for document-level MIAs, we find the MIA AUC to barely perform any better than random. We provide the results in Table VIII with the detailed experimental setup in Appendix B.
TABLE IV: Benchmarking MIAs on the MIMIR dataset [27] and Pythia-dedup-6.9B [6]. Sequences are sampled from the train (members) and test (non-members) split of The Pile [31]. MIA AUC computed using 500 members and 500 non-members.
@@
Limitations. While randomized, held-out test splits (when available) generally provide a proper fixed-model privacy game, the substantial overlap between members and non-members remains a core challenge. Post-hoc deduplication could potentially mitigate the overlap, yet determining the appropriate level of deduplication is difficult to evaluate and might distort the randomization. Its effectiveness likely varies depending on the dataset type, as demonstrated in the case of the GitHub subset. Utilizing the bag of words classifier to guide the appropriate level of deduplication could help achieve the right trade-off in practice.
@@
The practical significance of the overlap issue is a nuanced, perhaps philosophical, question. To rigorously examine what an LLM memorizes and how effectively MIAs can distinguish between members and non-members, one might prefer a setup with minimal overlap. Yet, overlap remains an inherent feature of text and it might be appropriate to account for overlap when inferring membership in real-world applications. For instance, when assessing whether a copyright-protected book was used to train an LLM, well-known quotes from it may be widespread online and included as near-duplicates in the training data. To identify whether the book was used in its entirety to train the model, it is likely essential to also account for such overlap. Whether MIAs, to e.g. study privacy risk, or detect copyright violations in practice, should be evaluated on datasets with an overlap thus remains an open question.
VI-B Injection of randomized unique sequences
To study sequence-level memorization, instead of relying on a randomized test split of documents, prior work has also proposed the injection of highly unique sequences.
For smaller language models, the controlled injection of specific sequences in the training data has been studied extensively. Henderson et al. [36] injects a variety of key-value pairs in a sequence-to-sequence dialogue model, while Carlini et al. [14] proposes the injection of canaries, unique sequences of random numbers, in the training set of LSTM models; both to measure how these sequences are memorized. Further work has, for instance, injected canaries to study memorization of federated learning [91] or of word embeddings [92].
More recently, Wei et al. [99] investigates how data watermarks, i.e. random sequences of tokens injected into the training dataset, are memorized by small scale language models (70M parameters trained on 100M tokens). Further, Meeus et al. [69] injected canaries (called traps sequences) into the pretraining dataset of CroissantLLM [30], a 1.3B parameter model trained on 3T tokens. The canaries are synthetically generated, unique sequences, injected with varying sequence length, number of repetitions, and perplexity. As the canaries are synthetically generated, by sampling from a pretrained LLM, the same process allows generating new, unique canaries coming from the exact same distribution, yet not part of the training dataset of CroissantLLM [30] (non-members). Lastly, Zhang et al. [108] also states that inferring membership of specially crafted canaries offers a viable path for proving whether LLMs are trained on certain data.
While Meeus et al. [69] use this setup to study how the inclusion of canaries enables content owners to track whether their content has been used for LLM training, the setup also provides a randomized and controlled setup to evaluate sequence-level MIAs against LLMs. We confirm this to be the case by fitting the bag-of-words classifier on copyright traps (using $500$ members and non-members), recovering an AUC of $0.543\pm 0.037$ for sequences of $100$ tokens.
Compared to the setup available with a randomized test split (Sec. VI-A), the use of randomized unique sequences to evaluate MIAs might have distinct benefits.
First, this setup is designed for sequences and eliminates the need to sample sequences from train and test splits.
Further, when the randomized sequences are highly unique, they could effectively eliminate the challenge of overlap between members and non-members. Meeus et al. [69] generate synthetic sequences of high perplexity, suggesting that these sequences would likely have less overlap with the rest of the target model training data than a randomized test split.
Lastly, injected sequences can be repeated. As it is well established that repeated sequences are memorized more [12, 44], injecting sequences with a varying number of repetitions would provide an MIA evaluation setup across multiple regimes of LLM memorization. This could enable the comparison of different MIAs with (more) statistical significance.
We instantiate MIAs from the literature on the dataset of member and non-member canaries as made available by Meeus et al. [69] on the open-source CroissantLLM [30]. Table V summarizes the MIA AUC across a number of repetitions. For $1,000$ repetitions, we recover MIAs reaching a performance significantly better than a random guess, with Ratio-LLaMA-2 [15] (AUC of $0.70$) and Min-KProb [85] (AUC of $0.63$) being the best performing methods.
TABLE V: Benchmarking MIAs across copyright traps for CroissantLLM [30]: MIA AUC for $500$ members and non-members.
@@ VI-B Injection of randomized unique sequences
Limitations. First, the MIA performance might only represent a worst-case risk and be not representative of real-world data. Indeed, highly unique sequences mitigate any overlap between members and non-members, making it easier for MIAs to succeed, especially when repeated during training. Moreover, Meeus et al. [69] shows that sequences with larger perplexity are more vulnerable, and considers sequences with a perplexity higher than naturally occurring ones. Hence, evaluating MIAs on unique sequences might give a worst-case MIA risk compared to real-world data. While this might be beneficial for some use-cases (e.g. privacy auditing of methods with formal guarantees [61, 62, 90], detecting copyright-protected content [69, 99]), evaluating MIAs on (artificially) unique sequences might not be representative of risks in practice. Notably, MIAs are likely to advance over time, and focusing on higher-risk sequences could establish a safety margin.
@@
Second, the success of MIAs might depend on certain properties of the sequence. For instance, the Ratio MIA [15] relies on the use of a reference model, which should normalize the membership score by how ‘complex’ a certain sequence is compared to other natural language. Surely, the extent to which this normalization is effective for an MIA against a target LLM depends on the choice of reference model, its training data and how well it generalizes to unseen data. Meeus et al. [69] use LLaMA-2 7B [94] to synthetically generate the trap sequences, likely affecting the success of certain reference models in detecting such traps. Also Duan et al. [15] finds ‘choosing the right reference model for a target LLM challenging and largely empirical’. Moreover, we hypothesize that certain MIAs may vary in effectiveness based on sequence length (e.g. Min-KProb [85], which relies on a subset of token-level probabilities). Similarly, some MIAs (e.g. SMIA [72]) depend on the semantic meaning of a target sequence, which might be less effective for sequences with very high perplexity due to the lack of semantically meaningful embeddings in such cases. Using randomized unique sequences as an evaluation setup could thus influence the relative performance of MIAs.
VI-C Finetuning
@@
While most of the prior work discussed in this paper consider pretrained LLMs as the target model, we believe that leveraging finetuning for MIA evaluation merits a discussion as part of the potential solutions. In particular, we find that new MIA methods proposed in recent work are reported to outperform prior methods [109, 100, 104, 107, 96, 72], while they are all evaluated on flawed setups using pretrained LLMs. Leveraging finetuning for MIA evaluation could enable a computationally feasible way for more reliable benchmarking. We further identify two additional benefits.
First, similar to the injection of unique sequences with varying repetitions, finetuning could be done for varying memorization regimes (adjusting e.g. number of parameters updated during finetuning, dataset size, learning rate [70]). As such, this setup would be well suited to compare MIA performance across methods.
Further, from a practical perspective, it is very sensible to study MIAs in posttraining attack models. Finetuning (or continued pretraining) is often done on smaller, domain-specific, high quality datasets which are often more susceptible to contain sensitive information. It is cost effective to start off from generalist model checkpoints (e.g. LLaMA-3 [2], Mistral-7B [43]) and train on a domain-specific, causal language modeling task for a few billions of tokens to specialize models. Examples include finetuning for legal domain adaptation [19], for multilingual abilities [3], and for medical domain adapation [49]. Moreover, annealing, or post-training techniques in which LLMs are trained on high-quality data splits after their initial pretraining phase, is now understood to be essential to boost model performance [38, 2]. Such high-quality data (news articles, textbooks, synthetic data) is often the category most likely to be copyright-protected.
Limitations. For some use-cases, including detecting the use of copyright-protected content or evaluation benchmarks during pretraining, the evaluation of MIAs remains desired for pretrained rather than for finetuned models. How MIA evaluation on finetuned models translates to larger models, datasets or other hyperparameters remains an open question [5].
VI-D Post-hoc control methods